OpenStudy (anonymous):

Kara's family is installing a circular pool in their back yard. The pool has a diameter of 12 feet and a circumference of 38.0 feet. What are the dimensions of Kara's pool, in meters (m)?

OpenStudy (anonymous):

By dimensions is the question asking for the box that fits perfectly around the circle?

OpenStudy (anonymous):

no

hero (hero):

Convert feet to meters

OpenStudy (nottim):

perimeter, area, volume? well, perimeter eh, since meters. but then circumference is there already

OpenStudy (nottim):

oh. duhhh

OpenStudy (anonymous):

A. diameter: 1.8 m, circumference: 11.6 m B. diameter: 3.7 m, circumference: 11.6 m C. diameter: 4.0 m, circumference: 12.7 m D. diameter: 4.7 m, circumference: 15.0 m these are my options but i dont get to any of them? what do i do?

hero (hero):

I gave you a hint. Do you know how to convert feet to meters?

OpenStudy (nottim):

should i say the conversion, hero?

hero (hero):

Yes, that would be the 2nd Hint

OpenStudy (nottim):

plop

hero (hero):

NotTim, that wasn't exactly what I had in mind

OpenStudy (nottim):

ehh...no?

OpenStudy (anonymous):

I don't know how many feet are in a meter, but I know that there's around 2.54 cm in an inch, 12 inches in a foot, and 100 cm in a meter. Dimensional analysis: Anything over itself is equal to 1. a/a = 1 2ft/2ft = 1 etc... So, 1in/2.54 cm = 1 = 2.54cm/1in Also, anything multiplied by 1 is equal to itself. 1*a = a 2ft/2ft * 10mi = 10mi etc... Using this, we can strategically multiply the number we're trying to convert, 12ft, in such a way that eliminates the current units and leaves us with the one we desire. 12ft/1 * 1 = 12ft/1 12ft/1 * 12in/1ft = 12ft/1 Feet on top and bottom, cancels out. 12*12in = 12ft/1 144in/1 = 12ft/1 144in/1 * 2.54cm/1in = 12ft/1 Inches cancel out, multiply. 365.76cm/1 = 12ft/1 365.76cm/1 * m/100cm = 12ft/1 Centimeters cancel out, divide. 3.6576m = 12ft Rounds to 3.7m. The circumference of a circle is equal to 2πr = πd (r is radius, d is diameter which is twice the radius). This is just something you memorize. C = π * d = π * 3.7m C = 3.7πm C approx = 11.624m Answer is B. diameter: 3.7 m, circumference: 11.6 m

OpenStudy (nottim):

serrr-fump. copy, paste, and answa!

OpenStudy (anonymous):

thank you guys ill let you know if i get it right in a few! :p

OpenStudy (anonymous):

NotTim, I wrote all of that myself.

OpenStudy (nottim):

whoaaa.......dedication! congrats, piggie

hero (hero):

@bluepig148 , posting the conversion rate would have sufficed.

OpenStudy (anonymous):

i got it right thank u guys gtg peace! :)

OpenStudy (anonymous):

I value intuition more than memorization. I don't know the conversion rate. If this were a real situation or one similar to it, knowledge of dimensional analysis would have saved me - not a conversion rate once learned and then forgotten.

hero (hero):

It looks like you went around in one gigantic circle just to get to a result that was simple to compute.

OpenStudy (nottim):

KIS-Keep is simple, friend

OpenStudy (anonymous):

When I was learning about conversions every lesson and example problem was an exercise in dimensional analysis. I don't see why this question would have been asked if it wasn't about that. Also, quit nit picking at everything I say.

hero (hero):

I wasn't trying to nit-pick. I was just giving a suggestion

OpenStudy (nottim):

occam's razor!

hero (hero):

The more information you give, the more likely it is to confuse the poster.

OpenStudy (anonymous):

NotTim, I think understanding is simpler than memorization.

hero (hero):

Giving simple to follow step by step hints is more effective than posting all of that.

OpenStudy (nottim):

eh, simple is different for every1, eh?